OpenStudy (anonymous):

Find the exponential function y = Ce^kt that passes through the two given points (0,4),(5,1/2)

OpenStudy (dumbcow):

plug points into equation ... solve system of 2 equ for C and k

OpenStudy (dumbcow):

\[4 = C e^{0k}\] \[5 = C e^{k/2}\] e^0 = 1 , therefore C = 4 \[\rightarrow 5 = 4e^{k/2}\]

OpenStudy (dumbcow):

can you solve for k?

OpenStudy (anonymous):

i have already set it up like you did but got confuse, i solve for k but still get the answer wrong

OpenStudy (dumbcow):

what did you get for k?

OpenStudy (dumbcow):

it may be correct but they want it entered in different form

OpenStudy (dumbcow):

\[e^{\ln x} = x\] so since k is log of something, the log will be cancelled out because it is power of "e"

OpenStudy (anonymous):

thanks

OpenStudy (dumbcow):

yw anyway \[y = 4 (\frac{25}{16})^{t}\]